VISTA IOT MINUTES, 2010-04-23 15-17h Garching / 14-16h UK / 11-13h Chile (prepared by W. Hummel) * * * Present: Garching: Mark Neeser (MNe), Michael Hilker (MHi), Wolfgang Hummel (WHu), Lander de Bilbao (LdB), Santiago: Valentin D. Ivanov (VDI), Alain Smette (ASm) Paranal: Steffen Mieske (SMi) UK (via phone): Mike Irwin (MIw) Jim Emerson (JEm) * * * Agenda ====== 1. VISTA status and planned activities (TSz) (intervention 24.04-5.05.2010) 2. VISTA Relevant PPRS Summary (VDI) and operational issues, including: - keyword CRVAL1,2=0 issue - night reports not available in DB 3. Progress of the surveys: -- Execution times & survey completion (AI 2009-01) -- Comparison of the current progress with the baseline simulation (AI 2009-14, AI 2010-03) 4. Photometric zeropoints from science frames (several related tickets; WHu, VDI) 5. QC monitoring scoring of opslogs values (finally operational!) 6. Re-processing of VISTA data by CASU (JLe) 7. Feedback from the PIs 8. Action Items review 9. Others 10. Date of the next meeting (May 21 or 28, 2010) * * * 1. VISTA status and planned activities VDI confirms the schedule for the VISTA intervention. Last observing night is 2010-04-23. The warming up of the instrument will start on 2010-04-24. The main issues that will be addressed are the cooling lines and the shimming of the instrument. VDI: Some test frames with the combined filter NB980/NB990 have been taken. According to A Kaufer, the 4 GTO nights which use this special filter have to be approved by the DG. These filters have not been tested yet on the sky. Preliminary approval had been given by the previous DG, but is pending approval by current DG. JEm: The filters are in two different paths of the tray and the pipeline does not know the allocation between filter wavelength and detector. But the association is documented. The filter consists of 16 elements, one half of which are NB980 and the other half is NB990. In order for the pipeline processing to work, a keyword must be assigned a filter ID for each detector. MIw: The arrangement of the NB980,990 filters in the tray doesn't directly affect the pipeline in terms of QC and calibration since all the usual processing steps will work without knowing the exact arrangement. The only part of the pipeline that is sensitive to filter type is the photometric calibration, but here their colour equations, and hence calibration using 2mass, will be negligibly different and hence no problem either. The main problem will occur later when trying to do science with the data when the user will need to know, say from the FITS headers, which filter is which. SMi: Flats with this filter have been taken on 2010-04-22. All sixteen fits header extensions show the same INS.FILT1.NAME value. VDI: details will be discussed off-line AI: JEm will provide information on the distribution of the individual filters within the tray after confirming with the PI for (Matt Jarvis), and the camera scientist (Gavin Dalton). JEm asked about the status of the M1, which shows two patches of "yellowing" VDI: The recent zeropoints are not affected. MIw: The long term trend in the zero-points shows a decrease of average zero-point of ~7% over the period 15th October to end of February. This is not necessarily purely a sensitivity change but could also reflect seasonal variations. We will continue to monitor this and post updates on the technical photometry pages here. VDI: There is no lab equipment on Paranal to measure the reflectivity of M1. 2. VISTA Relevant PPRS Summary PPRS 35093 VDI: When the TCS searches for a guide star, it nevertheless continues if the guide star search failed, without a striking warning. JEm: The progress of the template was implemented by intention, but an additional alarm is preferable. VDI: end of May would be a good deadline for implementing this. PPRS 34222 detector #14 channel #11 issue can be closed PPRS detector #16 The unreliability of detector #16 is a known issue. There is nothing that can be done to fix this. The detector #16 problem should be properly documented, (preferably in the User Manual so that the user is informed) The related PPRS can be closed PPRS 35169 CRVAL1,2=0 issue VDI: According to DBCM, 20000 VIRCAM frames in the archive are affected. JEm/MIw: The problem of missing CRVAL1,2 keywords in the extension headers might also be related to incorrect rotation matrix header updates in dither sequences. VDI: VISTA operations has to be logged for several nights in more details to find the origin of the CRVAL1,2 failure. It followed a discussion on the CRVAL1,2 value issue. The contact person for database content management (DBCM) issues, including the repair of ingested frames is Adam Dobrzycki. The CRPIX1,2 values are different from chip to chip. They contain the WCS reference pixel, which is a virtual pixel in the center of the FOV. The CRVAL1,2 values are the same in all extensions. They are simply a copy of the RA DEC values. The keyword repair does not require any coordinate transformations. AI: write email to P.Baksai or S.Sandrock to log VISTA operations PPRS missing night logs in DB VDI: PSO is aware if this. The project is delayed due to more urgent issues that came up in frame of the earthquake in 2010-02. MHi: The night reports distributed by email are stored by USD 3. Progress of the surveys: SMi: Given that VISTA will be operated TIOs (there will be no longer scheduled VISTA support astronomers in the near future) the manual overwriting of priorities offered by the OT is a work-around. Instead the survey progress/lag should be taken into account by the OT, that overwriting becomes obsolete. summary of a longer discussion is: The UltraVista public survey progress is behind the expectations derived from the simulation. The reason for this is: The assumption, that the highest priority public surveys require the best weather is no longer true as the survey management plans have been iterated and observing constraints have partly been relaxed. PSO continues to overwrite the priorities offered by the OT as often as possible to compensate for the lag of the UltraVista survey. It was anyway agreed at the previous IOT that the issue will be monitored further, since potential countermeasures must be based on a larger set of executed OBs. 4. Photometric zeropoints VDI: The photometric zeropoints, obtained from the pipeline on Paranal show a dependency on crowded fields and the histograms are different from detector to detector, and from filter to filter (both in asymmetry and in the width of scatter). Zp histograms appear reasonable for the J, H, and Ks bands, but are very broad in Y and Z. It is most probably related to data reduction issues and less to flat fielding. The plots of the report covert Oct 15 to April 7. MIw: CASU needs to generate the identical histogram plots in order to confirm if this is truly a pipeline issue. SMi: The zeropoints in dense fields are weaker. MIw: One first has to remove the 7% longterm trend AI: VDI: provide a comparison with ISAAC photometric data AI: ASm: A check in how far the Bolivian winter climate phenomenon impacts the photometric zeropoints. WHu: The zeropoints change also from pawprint to pawprint in a correlated manner. 5. QC monitoring scoring of opslogs values WHu: QC group scores now so-called opslog QC values generated by the pipeline on Paranal, thus the QC feedback delay has been shortened from 14 days to 1 hour. AI WHu: provide information on this to CASU 6. Re-processing of VISTA data by CASU MIw: Reprocessing of VISTA data from 15th October is finished and the PIs and SV teams have been informed. All incoming data is now processed with this new version. The latest data received and processed is 30th March 2010. 7. Feedback from the PIs MHi reported an overlap issue between VHS and VIKING areas 8. Action Items review See below 9. Others MHi: Some VVV OB's are affected by a high ellipticity. The VV OBs have short exposure times and are executed with AO priority = normal, meaning the AO cycle is less frequently updated. At the moment there is no rule when an OB should be repeated. As USG only gets reports for the spot checked science OBs, there are probably some OBs with undetected high ellipticity. WHu: processed science OBs are proposed to be repeated if at least 3-4 out of 16 detectors show an ellipticity > 0.2. The ellipticity is obtained from the pipeline running in Garching; de-blending option switched on. The elliptical PSF are all aligned. In most cases the ellipticity increasing towards one edge of the FOV exceeding the ellip = 0.2 thresholds detectors at the rim of the FOV. MIw pointed out that a large ellipticity combined with undersampling is especially critical and affects the image subtraction. The CASU conclusion is that ellipticity > 0.2 is the critical limit. AI: WHu It should be investigated, if the quartile instead of the average or median over the 16 ellipticity values can be used. WHu: The pipeline provides an option to de-blend sources in crowded fields. In the pipeline installed on Paranal this option is switched OFF by default. The Garching version has switched the option ON since some time with good experience. Should the pipeline on Paranal use the same de-blending=ON default option ? VDI: yes AI: LDB: use a new default value for the de-blending option 10. Date of the next meeting (May 21 or 28, 2010) May 28 2010 Progress report on previous action items: * AI 2009-01: Monitor execution times & survey completion status is ---------- being followed up by MAr, MRe. 2009-07-29: there is a first prototype 2009-10-06: no news (MRe) 2009-12-01: no news 2010-03-23: the progress was reported earlier (see above; MRe) let's have another review on the next IOT when we will have more OBs executed (MAr) 2010-04-23: no news STATUS: ONGOING * AI 2009-11: TSz should find out if the constraints sets could be carried with the OBs, to the OS and the fits header of the VIRCAM files. Otherwise it is not easy to propagate these data by interfaces. 2009-07-29: TSz iterating with R. Schmutzer, but DFI (T. Bierwirth) will also need to be involved. For surveys, given the small range of programs it is expected that there are not too many different OBs, and this might need to be followed in a different/manual way in the beginning - this could imply a major change for the tools that cannot be done in the last moment. 2010-03-23: the keywords have been implemented except for the moon; there is a ticket with details; some testing done by VDI, no errors found; keep ongoing, try to complete during my next shift end of Mar - early Apr. 2010-04-23: VDI reports one formula is missing STATUS: ONGOING * AI 2009-14: GHu to check that the seeing distribution function in the scheduler for VIRCAM takes into account the finite (0.51") nominal resolution of the instrument. 2010-03-23: GHu reported that it is not clear if TaToo will be used for scheduling time on VISTA; decided to keep this AI open until it becomes clear how VISTA time will be scheduled. 2010-04-23 no news STATUS: ONGOING * AI 2010-01: TSz the calibration database should be updated with new flats and master calibrations and how to do it regularly, as no procedure is currently in place. 2010-01-26: WHu should give Thomas Bierwirth a USB disk with the updated calibrations before Feb7, 2010. Then on try to define a procedure for a regular update of reference frames on Paranal 2010-03-23: done in Feb; the frequency of updates should be three months, the next one will be in May; SMi will talk to DA's to find out how to clean the database, i.e. removing the old calibrations with a script? Once this is done, we close this AI. 2010-04-23: given the intervention, it is not necessary to update the calibDB now with pre-intervention calibrations. STATUS: ONGOING * AI 2010-02: TSz: how many standard stars fields should be taken? Effort in February to take up to 3 fields. Request to implement a change of jitter pattern by JLe, from 4 to 5 jitter. Try to keep calibrations plan withing 1 hrs total execution time. 2010-01-26: JLe suggests to discuss issues with PIs, and provide feedbacks to TSz, especially in the ZY band calibrations in particular. 2010-03-23: MRe had a discussion with P. Lucas (VVV co-PI) about the frequency of the standards, the feedback is pending, we will keep this open until we hear from him. VDI points that we will overcalibrate at the beginning but at some point we won't need to bother anymore with standards at all because we will have the entire sky calibrated in the NIR (unless one needs better accuracy or in special cases of high extinction regions). A short discussion about feeding this back to the PIs - at the moment it is premature (as well as the information about the expected extra year necessary to complete the surveys) because we are still unsure ourselves. 2010-04-23: no news (MRe not attending) STATUS: ONGOING * * New Action Items from 2010-03-23: AI 2010-03: completion rate of the surveys to be compared with the new simulation of scheduling which should now include real submitted OBs. This comparison should be used to review the ranking and decide if it is necessary to revise it for the future. MRe to provide completion rate until May 1. GHu+VDI to run a new simulation with realistic OB database. Deadline for this AI is May IOT. 2010-04-23: no news AI 2010-04: TSz to modify the AO priority/validity 2010-04-33: no news (TSz not attending) AI 2010-05: JLe to incorporate the EXTNAME in the CASU products 2010-04-23: no news (JLe not attending) AI 2010-06: TSz+VDI to improve the TIO training, to prepare an operational manual 2010-04-23: done STATUS: ONGOING AI 2010-07: JEm to fix the SADT - 2010-04-23: JEm will fix the most critical items on the bug list, but does not guarantee to fix items that go beyond the original requirements. Some critical items are to be discussed, like the control of parameter values that might be manually edited by the user. The PA bug was identified and will be fixed. The new catalog issue on UCAC-3 isn't solved yet. The tool will be delivered by the end of next week. Ancillary material (documentation) will be distributed. AI 2010-08: VDI to take care of the new IP by May 2010-04-23: will be done STATUS: ONGOING New Action Items from this meeting 2010-04-23: AI 2010-09: photometric zeropoints as calculated by the pipeline. AI: WHu: It should be investigated, if the quartile of the ellipticity can be used as a threshold for OB rejection. AI: LdB: use a new default value for the de-blending option of the pipeline